OpenStudy (anonymous):

The perimeter of a rectangular shaped lawn is 52 feet and the area is 144 square feet. What are the dimensions of the lawn?

jigglypuff314 (jigglypuff314):

Perimeter of a rectangle = 2(width) + 2(length) Area of a rectangle = (length) * (width) so you can get 2W + 2L = 52 L*W = 144 do you know how to solve system of equations like this?

OpenStudy (anonymous):

yes thank you